OpenStudy (yamyam70):

What does this mean ? cos θ > 0

OpenStudy (anonymous):

what does this mean x > 3

OpenStudy (yamyam70):

x is greater than three

OpenStudy (yamyam70):

so , cos theta is greater than zero, what about it?

OpenStudy (anonymous):

90> θ>0 then cosθ>0

OpenStudy (yamyam70):

is it in Quadrant 1 ? @nonamenever

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what is the question exactly?

OpenStudy (ranga):

cos(x) > 0 in the interval (-pi/2, +pi/2) Since cos is a periodic function, you can add any integer multiple of 2pi to the above interval and cos will be > 0.

OpenStudy (yamyam70):

Find the exact value of the expression. sin θ = − 1/ 5 and cos θ > 0; find tan θ.

OpenStudy (ranga):

sin(x) = -1/5 = -0.2 x = -11.54 degrees cos(-11.54) = 0.98 tan(x) = sin(x) / cos(x) = -0.2 / 0.98 = ?

OpenStudy (yamyam70):

in the sample problem , the answer is a fraction in radian form. they did not solve for the angle

OpenStudy (yamyam70):

which I assume,

OpenStudy (ranga):

There are different ways of solving the same problem and arriving at the same answer. You can draw a right angle and arrive at the same answer: 

OpenStudy (ranga):

Here they are asking for tan(x) so the answer will not be in radians. It will be a number. If they ask for x (or theta) then the answer will be in radians or degrees.

OpenStudy (ranga):

Without solving for the angle this is how you can find tan(x) in the diagram. They have given you sin(x) = -1/5. Sin(x) is opposite / hypotenuse which I have drawn and I have put a negative sign denoting a negative angle x. The side marked with a question is: sqrt(5^2 - (-1)^2) = sqrt(24) tan(x) = opposite / adjacent = -1 / sqrt(24) = -sqrt(24)/24 = sqrt(6)/12

OpenStudy (yamyam70):

@ranga I salute you ! thanks!

OpenStudy (ranga):

you are welcome.